sandy011:

Simplify the expression: 2(5a-4a)+6a2(5a−4a)+6a

jkmiec26:

is it =2(5 etc. or is it 6a squared?

sandy011:

ah its 6a 2(5

sandy011:

i forgot the space

jkmiec26:

Okay so what you do first is distribute

sandy011:

okay

jkmiec26:

What do you get when you distribute?

jhonyy9:

like a first idea inside these parentheses there are 5a -4a yes ?

sandy011:

im so confused now

jhonyy9:

why ? you ve wrote there - no ?

sandy011:

imma just ask a friend for the answer

jhonyy9:

2(5a-4a)+6a 2(5a−4a)+6a is this in this way ?

jhonyy9:

so then 5a-4a = a 2a +6a = ?

jhonyy9:

@Laylalyssa

jkmiec26:

no you have to distribute first

jkmiec26:

you have to distribute everything

jhonyy9:

what you wan distribute ? inside parentheses are 5a -4a = a

jkmiec26:

2(5a-4a)+6a2(5a−4a)+6a 10a-8a+30a^3-24a^3 +6a

jkmiec26:

that's how u distribute

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jkmiec26 2(5a-4a)+6a here said that is space 2(5a−4a)+6a 10a-8a+30a^3-24a^3 +6a \(\color{#0cbb34}{\text{End of Quote}}\)